PLEASE CHECK ANSWERS I AM NOT SURE THEY ARE CORRECT.Exercise 6-7 Sandhill Co. reports the following for the month of June DateExplanationUnits Unit Cost Total Cost 121 345 202 2
Answer #1

a) Calculate weighted average cost per unit

Unit Unit Cost Total Cost
Beginning inventory 121 5 605
Purchase 1 345 6 2070
Purchase 2 202 7 1414
Total 668 4089

Weighted average unit cost = 4089/668 = 6.121 Per unit

Calculate cost of ending inventory and cost of goods sold

FIFO LIFO Average cost
Cost of ending inventory (202*7+48*6) = 1702 (121*5+129*6) = 1379 250*6.121 = 1530
Cost of goods sold 4089-1702 = 2387 4089-1379 = 2710 418*6.121 = 2559